OpenStudy (anonymous):

The discipline quotient varied inversely as the square of the number of unruly students. If the discipline quotient was 300 when the number of unruly students was totaled 5, what was the discipline quotient when the number of unruly students totaled 10? Use the equal ratio method please.

OpenStudy (kropot72):

Let x be the discipline quotient when the number of unruly students totaled 10. Then we can write the following equation: \[\frac{300}{\frac{1}{5^{2}}}=\frac{x}{\frac{1}{10^{2}}}\] Now you just need to solve for x.